OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + 2y – z = 7 4x – y + 3z = –2 2x + 2y – z = 9

OpenStudy (anonymous):

WEll. z = x + 2y - 7.......(1) From 1st equaton. y = 4x +3z + 2 From 2nd Euation. From (1) y = 4x + 3x + 6y - 21 + 2 => y = (19 - 4x)/5 Alo it means in terms of x z = x + 2/5(19-4x) Putting this value inthe thid equation, We get entir equation in x. Solve for it. Gt the answer.